# Self-Organization and Spectral Mechanism of Attractor Landscapes in High-Capacity Kernel Hopfield Networks

### v4

@@ 2.2 Learning Algorithms and Capacity Theory
An alternative path to high capacity is refining the learning rule. Gardner’s pioneering work established the theoretical capacity limit for patterns to be linearly separable [6]. The pseudo-inverse rule [10] and perceptron learning provide mechanisms to approach this limit by uncorrelating patterns. The introduction of kernel methods brought non-linear separation capabilities to this domain. Support Vector Machines (SVMs) have been applied to train recurrent networks, interpreting dynamics as iterative classification [11, 12]. Our previous work established KLR as a highly effective rule for this purpose [13, 5]. However, these studies typically evaluate performance based on static classification metrics (margin) or final retrieval capacity. They rarely investigate the dynamical mechanism; specifically, it remains unclear how the learned weight spectrum shapes the global attractor landscape during the transient recall phase, which is the primary focus of this study.
@@ 3.1 Kernel Logistic Regression Hopfield Network
We consider a network of $N$ bipolar neurons, $\bm{x}\in\{-1,1\}^{N}$, trained to store $P$ patterns $\{\bm{\xi}^{\mu}\}_{\mu=1}^{P}$. The KLR learning rule determines the dual variables $\alpha_{\mu i}$ for each neuron $i$ by minimizing a regularized logistic loss. The network state evolves synchronously based on the input potential:
@@ 3.3 Theoretical Framework: Spectral Representation
Matrix Representation of Weights: While kernel Hopfield networks do not store an explicit $N\times N$ weight matrix, the effective linear operator $\bm{W}$ in the high-dimensional feature space $\mathcal{H}$ governs the dynamics. Crucially, this operator is finite-rank and can be fully characterized by the stored patterns and the learned dual variables. Let $\bm{\Phi}=[\phi(\bm{\xi}^{1}),\dots,\phi(\bm{\xi}^{P})]$ be the feature matrix. The weight operator can be written as:
@@ 5.2 Empirical Validation of the Theory
$\text{StableRank}(\bm{\alpha})=\frac{\sum_{i=1}^{P}\sigma_{i}^{2}}{\sigma_{1}^{2}}.$ (10)
@@ A.2 Proof of Proposition 1 (Stability-Spectrum Link)
The energy function is defined as $E(\bm{s})=-\frac{1}{2}\langle\phi(\bm{s}),\bm{W}\phi(\bm{s})\rangle$. Let $\bm{s}^{*}$ be a fixed point. The Hessian matrix $H=\nabla^{2}E(\bm{s}^{*})$ is given by:
$H=J^{\top}\bm{W}J,$ (11)
where $J$ is the Jacobian matrix of the feature map $\phi(\bm{s})$ at $\bm{s}^{*}$.
@@ A.3 Theoretical Basis for Spectral Concentration (Proposition 2)
Akira TamamoriHe received his B.E., M.E., and D.E. degrees from Nagoya Institute of Technology, Nagoya, Japan, in 2008, 2010, 2014, respectively. From 2014 to 2016, he was a Research Assistant Professor at Institute of Statistical Mathematics, Tokyo, Japan. From 2016 to 2018, he was a Designated Assistant Professor at the Institute of Innovation for Future Society, Nagoya University, Japan. From 2018 to 2020, he was an lecturer at Aichi Institute of Technology, Japan. He has been an Associate Professor at the same institute. He is also a Visiting Assistant Professor at Speech and Language Processing Laboratory, Japan. He is a member of the Institute of Electronics, Information and Communication Engineers (IEICE), the Acoustical Society of Japan (ASJ), and Asia-Pacific Signal and Information Processing Association (APSIPA).

### v5

@@ 1 Introduction
The remainder of this paper is organized as follows. Section 2 reviews related work, positioning our study within the context of capacity theory and kernel associative memories. Section 3 introduces the network model and defines the geometric metrics used for our analysis. Section 4 presents the phenomenological results, detailing the discovery of the “Ridge of Optimization” and the associated “Force Antagonism.” Section 5 provides the theoretical core of this work, deriving the spectral mechanism of stability and validating the “Spectral Concentration” hypothesis. Finally, Section 6 discusses the broader implications of our findings, and Section 7 concludes the paper.
@@ 2.2 Learning Algorithms and Capacity Theory
An alternative path to high capacity involves refining the learning rule. Gardner’s pioneering work established the theoretical capacity limit for patterns to be linearly separable [6]. The pseudo-inverse rule [10] and perceptron learning provide mechanisms to approach this limit by uncorrelating patterns.
The introduction of kernel methods brought non-linear separation capabilities to this domain. Early works extended associative memory models to feature spaces, incorporating kernel functions directly into the energy Hamiltonian [11] or utilizing kernelized pseudo-inverse rules [12]. Building on these foundations, recent approaches have applied discriminative learning, such as Support Vector Machines (SVMs) or Kernel Logistic Regression (KLR), to train recurrent networks by interpreting the dynamics as an iterative classification task. Our previous work established KLR as a highly effective rule for this purpose, achieving state-of-the-art capacity [13, 5].
However, these studies typically evaluate performance based on static classification metrics (e.g., margin) or final retrieval capacity. They rarely investigate the dynamical mechanism; specifically, it remains unclear how the learned weight spectrum shapes the global attractor landscape during the transient recall phase, which is the primary focus of this study.
@@ 3.1 Kernel Logistic Regression Hopfield Network
We consider a network of $N$ bipolar neurons, $\bm{s}\in\{-1,1\}^{N}$, trained to store $P$ patterns $\{\bm{\xi}^{\mu}\}_{\mu=1}^{P}$. The KLR learning rule determines the dual variables $\alpha_{\mu i}$ for each neuron $i$ by minimizing a regularized logistic loss. The network state $\bm{s}$ evolves synchronously based on the input potential:
@@ 3.3 Theoretical Framework: Spectral Representation
Matrix Representation of Weights: While kernel Hopfield networks do not store an explicit $N\times N$ weight matrix, the effective linear operator $\bm{W}$ in the high-dimensional feature space $\mathcal{H}$ governs the dynamics. Crucially, this operator is finite-rank and can be fully characterized by the stored patterns and the learned dual variables. Let $\phi:\mathbb{R}^{N}\to\mathcal{H}$ be the feature mapping associated with the RBF kernel, which maps the input state into a high-dimensional Reproducing Kernel Hilbert Space (RKHS) $\mathcal{H}$. Let $\bm{\Phi}=[\phi(\bm{\xi}^{1}),\dots,\phi(\bm{\xi}^{P})]$ be the feature matrix collecting the mapped patterns. The weight operator can be written as:
@@ 5.2 Empirical Validation of the Theory
$\text{StableRank}(\bm{\alpha})\coloneqq\frac{\sum_{i=1}^{P}\sigma_{i}^{2}}{\sigma_{1}^{2}}.$ (10)
@@ A.2 Proof of Proposition 1 (Stability-Spectrum Link)
The energy function is defined as $E(\bm{s})=-\frac{1}{2}\langle\phi(\bm{s}),\bm{W}\phi(\bm{s})\rangle$. Let $\bm{s}^{*}$ be a fixed point. The Hessian matrix $H=\nabla^{2}E(\bm{s}^{*})$ is dominated by the first-order term (Gauss-Newton approximation):
$H\approx J^{\top}\bm{W}J,$ (11)
where we define the sign convention such that $H$ represents the magnitude of the curvature (positive definite around a stable attractor), ignoring the negative sign from the energy definition. $J$ is the Jacobian matrix of the feature map $\phi(\bm{s})$ at $\bm{s}^{*}$.
@@ A.3 Theoretical Basis for Spectral Concentration (Proposition 2)
Akira TamamoriHe received his B.E., M.E., and D.E. degrees from Nagoya Institute of Technology, Nagoya, Japan, in 2008, 2010, 2014, respectively. From 2014 to 2016, he was a Research Assistant Professor at Institute of Statistical Mathematics, Tokyo, Japan. From 2016 to 2018, he was a Designated Assistant Professor at the Institute of Innovation for Future Society, Nagoya University, Japan. From 2018 to 2020, he was an lecturer at Aichi Institute of Technology, Japan. He has been an Associate Professor at Aichi Institute of Technology, Japan. He is a member of the Institute of Electronics, Information and Communication Engineers (IEICE), the Acoustical Society of Japan (ASJ), and Asia-Pacific Signal and Information Processing Association (APSIPA).